EPS-HEP99 Abstract # 000 Parallel sessions: Plenary sessions:

Measurement of Jet Distributions in Charged–Current DIS at HERA

H1 Collaboration,

DESY, Notkestr.85, D-22607 Hamburg, Germany

Jet production in charged–current DIS at HERA is studied at the highest accessible values of Q^{2} using data collected with the H1 detector in 1994-1997. Differential (1+1) and (2+1)-jet rates and related distributions are measured. The measured jet distributions are compared to QCD model predictions. Comparisons with the corresponding jet distributions in NC processes are made.